A body of mass 3kg and volume 4×10^-4m^3 is hung from balance graduated in newtons. What is the reading of the balance when the body is;

a. In air

b. Fully immersed in water

c. ⅓ immersed in water

d. Fully immersed in paraffin.(Density of water is 10^3kgn^-3 and Density of paraffin is 800kgm^-3


1
Expert's answer
2021-11-23T10:52:13-0500

a.



"W=mg=(3)(9.8)=29.4\ N"


b.



"F=mg-\rho gV=(3)(9.8)-(1000)(9.8)(0.0004)=25.5\ N"


c.



"F=mg-\frac{1}{3}\rho gV=(3)(9.8)-\frac{1}{3}(1000)(9.8)(0.0004)=25.2\ N"


d.



"F=mg-\rho_p gV=(3)(9.8)-(800)(9.8)(0.0004)=26.3\ N"